OpenStudy (summerd613):

What is 2a to the power of 3 + 2b the the power of 2. A=4 and b=2. WILL FAN AND MEDAL HELPPPPPPO!!!!

OpenStudy (kenialove):

So 2a^3+2b 2(4)^3+2(2) What's 3 + 2(2) = ?

OpenStudy (kenialove):

Correct! Ok now what's 2(4) = ?

OpenStudy (kenialove):

Great! So then what's 8^7 = ?

OpenStudy (kenialove):

Oh okay hold on

OpenStudy (kenialove):

So then it's this .. 2a^3+2b^2 2(4)^3+2(2)^2 2(4)^3+16 8^19 = ?